This christmas, your brother Jeff will be 2 years from being twice as old as his siter Megan. The sum of Jeff's age and 3 times Megan's age is 66. How old is Megan?
Answer by stanbon(75887)

This christmas, your brother Jeff will be 2 years from being twice as old as his siter Megan. The sum of Jeff's age and 3 times Megan's age is 66. How old is Megan?
-----------------
j = 2m-2
j + 3m = 66
----
Substitute for "j" and solve for "m":
2m-2 +3m = 66
5m = 68
m = 13 3/5 years Megan's age now)
-----------------
Solve for "j":
j = 2m-2
j = 2(68/5)-(10/5)
j = 126/6 = 25 1/6 years (Jeff's age now)
======================
Cheers,
Stan H.
======================